# Rendering Wireless Environments Useful for Gradient Estimators: A Zero-Order Stochastic Federated Learning Method

### v1

@@ Abstract
Federated learning (FL) is a novel approach to machine learning that allows multiple edge devices to collaboratively train a model without disclosing their raw data. However, several challenges hinder the practical implementation of this approach, especially when devices and the server communicate over wireless channels, as it suffers from communication and computation bottlenecks in this case. By utilizing a communication-efficient framework, we propose a novel zero-order (ZO) method with a one-point gradient estimator that harnesses the nature of the wireless communication channel without requiring the knowledge of the channel state coefficient. It is the first method that includes the wireless channel in the learning algorithm itself instead of wasting resources to analyze it and remove its impact. The two main difficulties of this work are that in FL, the objective function is usually not convex, which makes the extension of FL to ZO methods challenging, and that including the impact of wireless channels requires extra attention. However, we overcome these difficulties and comprehensively analyze the proposed zero-order federated learning (ZOFL) framework. We establish its convergence theoretically, and we prove a convergence rate of $O(\frac{1}{\sqrt[3]{K}})$ in the nonconvex setting. We further demonstrate the potential of our algorithm with experimental results, taking into account independent and identically distributed (IID) and non-IID device data distributions.
@@ I Introduction
ZO methods are a subfield of optimization that assume that first-order (FO) information or access to function gradients is unavailable. ZO optimization is based on estimating the gradient using function values queried at a certain number of points. The number of function queries depends on the assumptions of the problem. For example, in multi-point gradient estimates [1, 2], they construct the gradient by performing the difference of function values obtained at many random or predefined points. However, they assume that the stochastic setting stays the same during all these queries. For example, for functions $\theta\mapsto f(\theta,S)$ subject to a stochastic variable $S$, two-point gradient estimates have the form,
$g=d\frac{f(\theta+\gamma\Phi,S)-f(\theta-\gamma\Phi,S)}{2\gamma}\Phi,$
with $\theta\in\mathbb{R}^{d}$ the optimization variable, $\gamma>0$ a small value, and $\Phi$ a random vector with a symmetric distribution. By contrast, one-point estimates that use only one function value [3, 4, 5], principally obtained at a random point,
@@
assume that the settings are continuously changing during optimization. This is an important property as it resonates with many realistic applications, like when the optimization is performed in wireless environments or is based on previous simulation results. Recently, an appeal to ZO optimization is emerging in the machine-learning community, where optimizers are based on gradient methods. Examples include reinforcement learning [6, 7], generating contrastive explanations for black-box classification models [8], and effecting adversarial perturbations on such models [9, 10].
On the other hand, with the massive amounts of data generated or accessed by mobile devices, a growing research interest in both sectors of academia and industry [11] is focused on FL [12, 13], as it’s a practical solution for training models on such data without the need to log them to a server. A lot of effort has been invested in developing first-order [12, 14, 15] and second-order [16, 17] methods to improve the efficacy of FL. These methods typically require access to the gradient or the Hessian of the local objective functions in their implementation to solve the optimization problem. However, using and exchanging such information raises many challenges, such as expensive communication and computation and privacy concerns [18].
@@
There is more interest recently in learning over wireless environments [19, 20, 21, 22, 23, 24], with the increase of devices connected to servers through cellular networks. In this paper, we are interested in this scenario illustrated in Fig. 1. Similarly to the aforementioned work, we are examining the case of analog communications between the devices and the server. However, it is a challenging problem as when the information is sent over the wireless channel, it becomes subject to a perturbation induced by the channel. This perturbation is not limited to additive noise, as noise is, in fact, due to thermal changes at the receiver. The channel acts as a filter for the transmitted signal [25, 26],
@@
where $x$ and $\hat{x}\in\mathbb{R}^{d}$ are the sent and received signals, respectively. $H\in\mathbb{R}^{d\times d}$ is the channel matrix, and $n\in\mathbb{R}^{d}$ is the additive noise, both of which are stochastic, constantly changing, and unknown. In general, all these entities are considered to have complex values. However, as we are interested in sending only real scalar values and we are not interested in decoding without error the sent information (but rather on using a perturbing on loss function), the phase shift introduced by the channel is considered as an additional perturbation in the channel model. Since our model is based on perturbing the loss function, considering only real values in the equation above is sufficient. We elaborate further on the channel modeling and why we can consider it real in Appendix A for the interested reader. In FL, $x$ may denote the model or its gradients sent over the channel. To remove this impact, every channel element must be analyzed and removed to retrieve the sent information. This analysis is costly in computation and time resources. Thus, here our objective is to study FL in wireless environments without wasting resources.
Further, we are interested in exploring the potential of ZO optimization to deal with some of the difficulties demonstrated by FL. We then consider an FL setting where a central server coordinates with $N$ edge devices to solve an optimization problem collaboratively. The data is private to every device and the exchanges between the server and the devices is restricted to the optimization parameters. To that end, let $\mathcal{N}=\{1,...,N\}$ be the set of devices and $\theta\in\mathbb{R}^{d}$ denote the global model. We define $F_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ as the loss function associated with the local data stored on device $i$, $\forall i\in\mathcal{N}$. The objective is to minimize the function $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$ that is composed of the said devices’ loss functions, such that
@@ I-A Motivation for Our Work
Communication Bottleneck. In general, the main idea of FL is that the devices receive the model from the server, make use of their data to update the gradient, and then send back their gradients without ever disclosing their data. The server then updates the model using the collected and averaged gradients, and the process repeats. Since the gradients have the same dimension as the model, in every uplink step, there are $Nd$ values that need to be uploaded, which forms the fundamental communication bottleneck in FL. To deal with this issue, some propose local multiple gradient descent steps to be done by the devices before sending their gradients back to the server to save communication resources [27], or allow partial device participation at every iteration [28, 29], or both [12]. Others propose lossy compression of the gradient before uploading it to the server. For example, all [30, 31, 32] suggest stochastic unbiased quantization approaches, where gradients are approximated with a finite set of discrete values for efficiency. [33] proposes the quantization of gradient differences of the current and previous iterations, allowing the update to incorporate new information, while [34] proposes the sparsification of this difference. Sparsification means that if a vector component is not large enough, it will not be transmitted.
@@
Computation Demands. Unlike standard methods that rely on the computational capabilities of participating devices, our approach is less demanding. Devices simply receive the global model, query it with their data, and send back the scalar loss, eliminating the need for "backward pass" computation. Only the "forward pass" is performed.
Black-Box Optimization in FL. One motivation for employing ZO methods is black-box problems [35] when gradient information cannot be acquired or is complicated to compute. For example, in hyperparameter tuning, gradients cannot be calculated, as there isn’t an analytic relationship between the loss function and the hyperparameters [36].
I-B Challenges and Contribution
Addressing nonconvexity in FL is challenging. Our ZO method must handle nonconvexity, noise, and stochasticity efficiently, which can slow down convergence in gradient techniques. Additionally, the channel’s introduction adds uncertainty and constraints on the number of communication exchanges. We need to ensure consistent and reliable performance, considering unknown probability distributions and fewer function evaluations. Unlike convex cases, nonconvex optimization does not allow easy quantification of optimization progress. Verifying gradient convergence becomes intricate due to biased gradient estimates. Moreover, unbounded variance in one-point estimates can lead to significant gradient deviations. These challenges involve technical and intuitive complexities we navigate.
In this work, we overcome these difficulties and propose a new communication-efficient algorithm in the nonconvex setting. This algorithm differs from the standard gradient method as it entails two reception-update steps instead of one, and it is not a simple extension of FO to ZO where the devices still have to upload their full model/gradient, as is the case in [35]. By limiting the exchange to scalar-valued updates, we counter the communication bottleneck, and we save up to a factor of $O(d)$ per communication round, in comparison to standard methods, in terms of total exchanges of variables between the devices and the server, saving a lot of execution time. We harness the nature of autocorrelated channels for truly "blind" reception of the data. We prove the convergence theoretically with a one-point estimate and provide experimental evidence. An important distinction worth noting is that standard ZO methods establish convergence by focusing on the expected convergence of the exact gradient. In contrast to prior research, our approach goes further in the proof. We demonstrate the convergence of the exact gradient itself almost surely, not solely its expected value. The key element in this proof is employing Doob’s martingale inequality to constrain the stochastic error resulting from the estimated gradient. We finally extend the analysis to non-symmetrical channel models, i.e., channels without zero-mean, and thus provide a practical algorithm for general settings.
@@ II-A The 1P-ZOFL Algorithm
Input: Initial model $\theta_{0}\in\mathbb{R}^{d}$, the initial step-sizes $\alpha_{0}$ and $\gamma_{0}$, and the channels’ standard deviation $\sigma_{h}$
@@
3:     The server broadcasts $\theta_{k}+\gamma_{k}\Phi_{k}\sum_{j=1}^{N}\Big{(}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}$ to all devices
@@
The devices must carry out two communication steps. In the first, every device sends the value $\frac{1}{\sigma_{h}^{2}}$ to the server. According to equation (1), the server receives $\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}$ from every device $j$. Hence, it receives the sum in step $2$. Afterward, the server uses the values received to adjust the model and broadcasts it to the devices. When device $i$ receives the model, it receives $h_{i,k+1}^{DL}[\theta_{k}+\gamma_{k}\Phi_{k}\sum_{j=1}^{N}(\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k})]+n_{i,k+1}^{DL}$, and to simplify notation, we let the stochastic vector $[h_{i,k+1}^{DL},n_{i,k+1}^{DL}]$ be included within the big vector $S_{i,k+1}$ of stochastic perturbations. Device $i$ then queries this received model to obtain the stochastic loss $f_{i}$. Then the devices send $f$ to the server in the second communication step, and according to equation (1), the server receives the quantity indicated in step $4$. Finally, the server assembles the gradient estimate and is able to update $\theta$ according to step $7$. All transmissions are subject to channel scaling and additive noise. We designate them by $h$ and $n$ in the device-to-server transmission. In the server-to-devices one, we designate them by $S$. We let $\tilde{f}_{i}=\frac{f_{i}}{\sigma_{h}^{2}}$ be the normalized loss function and define $\alpha_{k}$ and $\gamma_{k}$ as two step-sizes and $\Phi_{k}\in\mathbb{R}^{d}$ as a perturbation vector generated by the server that has the same dimension as that of the model.
@@ II-B The Estimated Gradient
$n_{i,k}$ is assumed to be a zero-mean uncorrelated noise with bounded variance, meaning $E(n_{i,k})=0$ and $E(n_{i,k}^{2})=\sigma_{n}^{2}<\infty$, $\forall i\in\mathcal{N}$, $\forall k\in\mathbb{N}^{+}$. For any timeslot $k$, $E(n_{i,k}n_{j,k})=0$ if $i\neq j$. For any device $i$, $E(n_{i,k}n_{i,k^{\prime}})=0$ if $k\neq k^{\prime}$.
@@
Let $\Phi_{k}=(\phi_{k}^{1},\phi_{k}^{2},\ldots,\phi_{k}^{d})^{T}$. At each iteration $k$, the server generates its $\Phi_{k}$ vector independently from other iterations. In addition, the elements of $\Phi_{k}$ are assumed i.i.d with $\mathbb{E}(\phi_{k}^{d_{1}}\phi_{k}^{d_{2}})=0$ for $d_{1}\neq d_{2}$ and there exists $\alpha_{2}>0$ such that $\mathbb{E}(\phi_{k}^{d_{j}})^{2}=\alpha_{2}$, $\forall{d_{j}}$, $\forall k$. We further assume there exists a constant $\alpha_{3}>0$ where $\|\Phi_{k}\|\leq\alpha_{3}$, $\forall k$.
@@
An example of a perturbation vector satisfying Assumption 2, is picking every dimension of $\Phi_{k}$ from $\{-\frac{1}{\sqrt{d}},\frac{1}{\sqrt{d}}\}$ with equal probability. Then, $\alpha_{2}=\frac{1}{d}$ and $\alpha_{3}=1$.
@@
All loss functions $\theta\mapsto f_{i}(\theta,S_{i})$ are Lipschitz continuous with Lipschitz constant $L_{S_{i}}$, $|f_{i}(\theta,S_{i})-f_{i}(\theta^{\prime},S_{i})|\leq L_{S_{i}}\|\theta-\theta^{\prime}\|$, $\forall i\in\mathcal{N}$. In addition, $\mathbb{E}_{S_{i}}f_{i}(\theta,S_{i})<\infty,\forall i\in\mathcal{N}$.
@@
Let Assumptions 1 and 2 be satisfied and define the scalar value $c_{1}=\alpha_{2}\frac{K_{hh}}{\sigma_{h}^{4}}$, then the gradient estimator is biased w.r.t. the objective function’s exact gradient $\nabla F(\theta)$. Concretely, $\mathbb{E}[g_{k}|\mathcal{H}_{k}]=c_{1}\gamma_{k}(\nabla F(\theta_{k})+b_{k})$ $\forall k\in\mathbb{N}^{+}$, where $b_{k}$ is the bias term. Proof: Refer to Appendix B-A.
@@
Let Assumptions 1-3 and the inequality $\|\theta_{k}\|<\infty$ hold almost surely. There exist a bounded constant $c_{2}>0$, such that $\mathbb{E}[\|g_{k}\|^{2}|\mathcal{H}_{k}]\leq c_{2}$ almost surely. Proof: Refer to Appendix B-B.

### v2

@@ Abstract
Cross-device federated learning (FL) is a growing machine learning setting whereby multiple edge devices collaborate to train a model without disclosing their raw data. With the great number of mobile devices participating in more FL applications via the wireless environment, the practical implementation of these applications will be hindered due to the limited uplink capacity of devices, causing critical bottlenecks. In this work, we propose a novel doubly communication-efficient zero-order (ZO) method with a one-point gradient estimator that replaces communicating long vectors with scalar values and that harnesses the nature of the wireless communication channel, overcoming the need to know the channel state coefficient. It is the first method that includes the wireless channel in the learning algorithm itself instead of wasting resources to analyze it and remove its impact. We then offer a thorough analysis of the proposed zero-order federated learning (ZOFL) framework and prove that our method converges almost surely, which is a novel result in nonconvex ZO optimization. We further prove a convergence rate of $O(\frac{1}{\sqrt[3]{K}})$ in the nonconvex setting. We finally demonstrate the potential of our algorithm with experimental results.
@@ I Introduction
ZO methods are a subfield of optimization that assume that first-order (FO) information or access to function gradients is unavailable. This arises in scenarios where exact gradient calculation may be expensive or unfeasible due to the inaccessibility of closed-form objectives. ZO optimization is based on estimating the gradient using function values queried at a certain number of points. The number of function queries depends on the assumptions of the problem. For example, in multi-point gradient estimates [1, 2], they construct the gradient by performing the difference of function values obtained at many random or predefined points. However, they assume that the stochastic setting stays the same during all these queries. By contrast, one-point estimates that use only one function value [3, 4], principally obtained at a random point,
@@
assume that the settings are continuously changing during optimization. This is an important property as it resonates with many realistic applications, like when the optimization is performed in wireless environments or is based on previous simulation results. Recently, an appeal to ZO optimization is emerging in the machine-learning community, where optimizers are based on gradient methods. Examples include reinforcement learning [5], generating contrastive explanations for black-box classification models [6], and effecting adversarial perturbations on such models [7].
On the other hand, with the massive amounts of data generated or accessed by mobile devices, a growing research interest in both sectors of academia and industry is focused on FL [8, 9], as it is a practical solution for training models on such data without the need to log them to a server. A lot of effort has been invested in developing first-order [9, 10, 11] and second-order methods [12, 13] to improve the efficacy of FL. These methods typically require access to the gradient or the Hessian of the local objective functions in their implementation to solve the optimization problem. However, using and exchanging such information raises many challenges, such as expensive communication and computation and privacy concerns [14].
@@
There is more interest recently in learning over wireless environments [15, 16, 17, 18], with the increase of devices connected to servers through cellular networks. In this paper, we are interested in this scenario illustrated in Fig. 1. Similarly to the aforementioned work, we are examining the case of analog communications between the devices and the server. However, it is a challenging problem as when the information is sent over the wireless channel, it becomes subject to a perturbation induced by the channel. This perturbation is not limited to additive noise, as noise is, in fact, due to thermal changes at the receiver. The channel acts as a filter for the transmitted signal [19],
@@
where $x$ and $\hat{x}\in\mathbb{C}^{d}$ are the sent and received signals, respectively. $H\in\mathbb{C}^{d\times d}$ is the channel matrix, and $n\in\mathbb{C}^{d}$ is the additive noise, both of which are stochastic, constantly changing, and unknown. In general, all these entities are considered to have complex values. However, as we are interested in sending only real scalar values and we are not interested in decoding without error the sent information (but rather in perturbing the loss function), the phase shift introduced by the channel is considered as an additional perturbation in the channel model. Since our model is based on perturbing the loss function, considering only the real part in the equation above is sufficient. We elaborate further on the channel modeling in Appendix A for the interested reader. In FL, $x$ may denote the model or its gradients sent over the channel. To remove this impact, every channel element must be analyzed and removed to retrieve the sent information. This analysis is costly in computation and time resources. Thus, here our objective is to study FL in wireless environments without wasting resources. Further, we are interested in exploring the potential of ZO optimization to deal with some of the difficulties demonstrated by FL.
We thus consider an FL setting where a central server coordinates with $N$ edge devices to solve an optimization problem collaboratively. The data is private to every device and the exchanges between the server and the devices is restricted to the optimization parameters. To that end, let $\mathcal{N}=\{1,...,N\}$ be the set of devices and $\theta\in\mathbb{R}^{d}$ denote the global model. We define $F_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ as the loss function associated with the local data stored on device $i$, $\forall i\in\mathcal{N}$. The objective is to minimize the function $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$ that is composed of the said devices’ loss functions, such that
@@ I-A Motivation for Our Work
Communication Bottleneck. In general, the main idea of FL is that the devices receive the model from the server, make use of their data to update the gradient, and then send back their gradients without ever disclosing their data. The server then updates the model using the collected and averaged gradients, and the process repeats. Since the gradients have the same dimension as the model, in every uplink step, there are $Nd$ values that need to be uploaded, which forms the fundamental communication bottleneck in FL. To deal with this issue, some propose local multiple gradient descent steps to be done by the devices before sending their gradients back to the server to save communication resources [21], or allow partial device participation at every iteration [22], or both [9]. Others propose lossy compression of the gradient before uploading it to the server. For example, [23] suggests a stochastic unbiased quantization approach where gradients are approximated with a finite set of discrete values. [24] proposes the quantization of gradient differences between the current and previous iterations, allowing the update to incorporate new information, while [25] proposes the sparsification of this difference.
@@
I-B Contribution
In this work, we propose a new communication-efficient algorithm in the nonconvex setting. This algorithm differs from the standard gradient method as it entails two reception-update steps instead of one, and it is not a simple extension of FO to ZO where the devices still have to upload their full model/gradient, as is the case in [26]. By limiting the exchange to scalar-valued updates, we counter the communication bottleneck, and we save up to a factor of $O(d)$ per communication round, in comparison to standard methods, in terms of total exchanges of variables between the devices and the server, saving a lot of execution time. We harness the nature of autocorrelated channels for truly "blind" reception of the data. We prove the convergence theoretically with a one-point estimate and provide experimental evidence. An important distinction worth noting is that standard ZO methods establish convergence by focusing on the expected convergence of the exact gradient. In contrast to prior research, our approach goes further in the proof. We demonstrate the convergence of the exact gradient itself almost surely, not solely its expected value. The key element in this proof is employing Doob’s martingale inequality to constrain the stochastic error resulting from the estimated gradient. We further demonstrate a convergence rate of $O(\frac{1}{\sqrt[3]{K}})$ which outmatches centralized algorithms in the nonconvex ZO setting with one-point or noisy two-point estimates [27, 28]. While this rate matches our previous work [29], the latter explores a ZO-distributed method aiming at achieving consensus. It entails a gradient tracking technique, which is a distinct algorithm that involves sharing between neighbors and updating two variables: the optimization variable and an auxiliary one. Thus, it has a worse communication cost as each user must send a long vector of $2d$ values, whereas here, only $2$ scalar values are sent per user. Furthermore, contrary to this work, the gradient tracking algorithm’s performance is dependent on the network architecture, and the channel is not included in the algorithm. This simplifies the analysis regarding the gradient estimate and its bias as they’re independent of the channel variables and the intertwined stochastic associations between them.
@@ II-A The 1P-ZOFL Algorithm
Input: $\theta_{0}$, $\alpha_{0}$, $\gamma_{0}$, $\sigma_{h}$
@@
3:     The server broadcasts $\theta_{k}^{\prime}$ to all devices, where $\theta_{k}^{\prime}=\theta_{k}+\gamma_{k}\Phi_{k}\sum_{j=1}^{N}\big{(}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\big{)}$
@@
The devices must carry out two communication steps. In the first, every device sends the value $\frac{1}{\sigma_{h}^{2}}$ (or any other scalar value) to the server. According to equation (1), the server receives $\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}$ from every device $j$. Hence, it receives the sum in step $2$. Afterward, the server uses the values received to adjust the model and broadcasts it to the devices. When device $i$ receives the model, it receives $h_{i,k+1}^{DL}[\theta_{k}+\gamma_{k}\Phi_{k}\sum_{j=1}^{N}(\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k})]+n_{i,k+1}^{DL}$, and to simplify notation, we let the stochastic vector $[h_{i,k+1}^{DL},n_{i,k+1}^{DL}]$ be included within the big vector $S_{i,k+1}$ of stochastic perturbations. Device $i$ then queries this received model to obtain the stochastic loss $f_{i}$. Then the devices send $f$ to the server in the second communication step, and according to equation (1), the server receives the quantity indicated in step $4$. Finally, the server assembles the gradient estimate and is able to update $\theta$ according to step $7$. All transmissions are subject to channel scaling and additive noise. We designate them by $h$ and $n$ in the device-to-server transmission. In the server-to-devices one, we designate them by $S$. We let $\tilde{f}_{i}=\frac{f_{i}}{\sigma_{h}^{2}}$ be the normalized loss function and define $\alpha_{k}$ and $\gamma_{k}$ as two step-sizes and $\Phi_{k}\in\mathbb{R}^{d}$ as a perturbation vector generated by the server that has the same dimension as that of the model.
@@ II-B The Estimated Gradient
$n_{i,k}$ is assumed to be a zero-mean uncorrelated noise with bounded variance, meaning $E(n_{i,k})=0$ and $E(n_{i,k}^{2})=\sigma_{n}^{2}<\infty$, $\forall i\in\mathcal{N}$, $\forall k\in\mathbb{N}^{+}$.
@@
Let $\Phi_{k}=(\phi_{k}^{1},\phi_{k}^{2},\ldots,\phi_{k}^{d})^{T}$. At each iteration $k$, the server generates its $\Phi_{k}$ vector independently from other iterations. In addition, the elements of $\Phi_{k}$ are assumed i.i.d with $\mathbb{E}(\phi_{k}^{d_{1}}\phi_{k}^{d_{2}})=0$ for $d_{1}\neq d_{2}$ and there exists $\beta_{1}>0$ such that $\mathbb{E}(\phi_{k}^{d_{j}})^{2}=\beta_{1}$, $\forall{d_{j}}$, $\forall k$. We further assume there exists a constant $\beta_{2}>0$ where $\|\Phi_{k}\|\leq\beta_{2}$, $\forall k$.
@@
Generating $\Phi_{k}$ such that $\mathbb{P}(\phi_{k}^{d_{j}}=-\frac{1}{\sqrt{d}})=\mathbb{P}(\phi_{k}^{d_{j}}=\frac{1}{\sqrt{d}})=\frac{1}{2}$, $\forall{d_{j}}$, $\forall k$. Then, $\beta_{1}=\frac{1}{d}$ and $\beta_{2}=1$.
@@
All loss functions $\theta\mapsto f_{i}(\theta,S_{i})$ are Lipschitz continuous with Lipschitz constant $L_{S_{i}}$, $|f_{i}(\theta,S_{i})-f_{i}(\theta^{\prime},S_{i})|\leq L_{S_{i}}\|\theta-\theta^{\prime}\|$, $\forall i\in\mathcal{N}$. In addition, we assume there exists $C>0$ such that $|f_{i}(\theta,S_{i})|\leq C<\infty,\forall i\in\mathcal{N}$.
@@
Let Assumptions 1 and 2 be satisfied and define the scalar value $c_{1}=\beta_{1}\frac{K_{hh}}{\sigma_{h}^{4}}$, then the gradient estimator is biased w.r.t. the objective function’s exact gradient $\nabla F(\theta)$. Concretely, $\mathbb{E}[g_{k}|\mathcal{H}_{k}]=c_{1}\gamma_{k}(\nabla F(\theta_{k})+b_{k})$, $\forall k\in\mathbb{N}^{+}$, where $b_{k}$ is the bias term.
Proof: Refer to Appendix B-A.
@@
Let Assumptions 1-3 hold. There exist a bounded constant $c_{2}>0$, such that $\mathbb{E}[\|g_{k}\|^{2}|\mathcal{H}_{k}]\leq c_{2}$.
Proof: Refer to Appendix B-B.